### v1

# Correlative and Discriminative Label Grouping for Multi-Label Visual Prompt Tuning

Abstract
Modeling label correlations has always played a pivotal role in multi-label image classification (MLC), attracting significant attention from researchers. However, recent studies have overemphasized co-occurrence relationships among labels, which can lead to overfitting risk on this overemphasis, resulting in suboptimal models. To tackle this problem, we advocate for balancing correlative and discriminative relationships among labels to mitigate the risk of overfitting and enhance model performance. To this end, we propose the Multi-Label Visual Prompt Tuning framework, a novel and parameter-efficient method that groups classes into multiple class subsets according to label co-occurrence and mutual exclusivity relationships, and then models them respectively to balance the two relationships. In this work, since each group contains multiple classes, multiple prompt tokens are adopted within Vision Transformer (ViT) to capture the correlation or discriminative label relationship within each group, and effectively learn correlation or discriminative representations for class subsets. On the other hand, each group contains multiple group-aware visual representations that may correspond to multiple classes, and the mixture of experts (MoE) model can cleverly assign them from the group-aware to the label-aware, adaptively obtaining label-aware representation, which is more conducive to classification. Experiments on multiple benchmark datasets show that our proposed approach achieves competitive results and outperforms SOTA methods on multiple pre-trained models.
1 Introduction
Multi-label image classification (MLC) aims to assign multiple labels to a single instance [56]. As a fundamental machine learning algorithm [31], MLC reflects real-world scenarios and is an essential component in many applications, e.g., face recognition [3], and scene understanding [57], weakly-supervised semantic segmentation [2]. Compared with single-label settings, the MLC presents significant challenges due to the exponential expansion of the output space, which usually requires large-scale datasets [45].
In real-world scenarios, co-occurrence between labels is prevalent, and numerous studies [47, 55] have leveraged label relationships to address the challenges resulting from complex label output spaces. In the context of deep learning, pioneering works [41, 44] employ recurrent neural networks (RNNs) to capture limited correlations between labels, while later studies [7, 46] utilize graph convolutional networks (GCNs) to capture global relationships in graphs or higher-order relationships in hypergraphs. More recently, for better representations, most researchers employ attention architectures to decouple an image into multiple label-aware features, as seen in methods like SSGRL [5], Q2L [30], etc.
(a) predicted image
(b) predicted probability
Figure 1: This illustrates the negative impact of co-occurrence. Due to the high co-occurrence probability $p(\text{{dog}}|\text{{frisbee}})$, e.g., 0.23, the model incorrectly infers the presence of dog when the label frisbee appears.
Despite the great advantages these methods showcase, two significant drawbacks require mitigation. The pre-trained model in these methods needs to be fine-tuned, leading to increased computational and resource demands. Also, for large-scale pre-trained models (e.g., DINOv2 [35]), full fine-tuning can negatively affect performance on downstream tasks due to object tasks and data distributions [25]. More importantly, while modeling correlations between labels can enhance model performance, overemphasizing these correlations will lead to the risks of overfitting and may misguide the model to make erroneous inferences, resulting in sub-optimal overall performance [28]. Figure 1 verifies this in an empirical example. Recent attempts to address this issue, such as BoostMLC [50] and PAT [49], still involve considerable computational costs (e.g., multiple teacher models and more augmented images).
Recently, visual prompt tuning (VPT) [21] has emerged as a parameter-efficient fine-tuning (PEFT) [11, 14] method for adapting large-scale pre-trained vision transformer models to downstream tasks [53, 43, 17]. Naturally, the success of VPT has inspired us to explore using PEFT to address the first issue. However, existing VPT-like methods primarily downstream tasks in single-label prediction [16, 36] or dense prediction tasks [22, 19], which may not suit MLC where the correlation between labels is not considered during the learning representation phase.
Regarding the second drawback, the key is to mitigate the adverse effects of overemphasizing co-occurrence relationships while preserving their advantages. A Naïve idea is to model the correlative (CO) and discriminative (DC) relationships between labels separately and then balance them. In terms of CO, a reasonable strategy is to highlight class pairs with high co-occurrence probabilities within a co-occurrence graph [7, 5], thereby modeling the CO between these classes. In contrast, DC tends to emphasize the characteristics of the object class to highlight distinguishability. Therefore, class pairs with low co-occurrence probability may have higher DC relationships (discriminability). Existing studies [50, 49] tend to be one of them or require more complex models and computation. Therefore, to effectively balance these two relationships, we can naturally group the classes into multiple sets according to their co-occurrence probabilities. All classes are placed together in the CO groups via high probabilities. Conversely, for DC groups, it is advocated to divide all classes using low co-occurrence probabilities.
Drawing on the above investigations and analysis, we propose a multi-label visual prompt tuning (ML-VPT) framework that utilizes grouping classes to balance the two relationships simply and effectively. Firstly, we take advantage of VPT and introduce an equal number of visual prompt tokens to each group, which facilitates learning visual representations for each class subset, referred to as group-aware representations. Unlike mainstream methods [5, 30], our approach offers two great advantages: It emphasizes directly modeling label relationships based on multiple prompt tokens for each group within the visual encoder. Our method maintains a balanced focus, not overemphasizing the CO relationship between labels while considering DC.  Secondly, each class subset comprises multiple classes, and each group incorporates multiple visual prompt tokens, resulting in each group containing multiple group-aware representations. To more effectively highlight the distinctions between classes, we employ a MoE model to selectively identify suitable visual representations for each class within the relevant group-aware representations. MoE can simplify the mapping from group-aware to label-aware complexity and enable adaptive mapping tailored to specific images.
Overall, our contribution is two-fold: 1) We propose a novel framework for MLC that groups classes first and subsequently models them within ViT. The proposed method balances the CO and DC relationships to overcome the risk of co-occurrence overfitting and improve model performance. 2) The group-aware MoEs are implemented to map group-aware representations to label-aware ones. In a dynamic way, multiple appropriate experts with gating networks are chosen to construct label-aware representations.
2 Related Work
Multi-Label Image Classification. MLC aims to predict multiple labels associated with an instance simultaneously [48]. The mainstream MLC methods can be roughly summarized into two types: modeling label correlations and learning label-aware representations. The former typically employs GCNs or RNNs to capture dependencies between labels [41, 44, 7, 5], while the latter often utilizes attention mechanisms to learn category-specific representations, which focus on effectively processing contextual information in images [51, 30, 38, 33, 34]. Very recently, limited studies have begun to investigate overfitting and causal interventions in the context of MLC. For example, CDD [28] and IDA [29] eliminate the contextual debiasing. BoostMLC [50] and PAT [49] attempt to eliminate the risk of overfitting resulting from overemphasizing co-occurrence, but involve complex calculation processes and computational overhead.
Visual Prompt Tuning. To strike an optimal balance between computational cost and performance, Jia et al. pioneer VPT [21], which effectively adapts transformer models pre-trained on large-scale datasets for various downstream tasks. Subsequently, research initiatives analogous to VPT emerged [53, 43, 17], with many researchers employing VPT across various computer vision applications, such as classification [36], test-time adaptation [16], segmentation [22], and continual learning [15]. However, these studies are unsuitable for MLC, as they fail to account for label correlations, whereas our approach explicitly considers them.
Mixture of Experts. The MoE [20] dynamically selects specialized expert networks to process input data, enhancing overall model performance. MoE has been successfully applied across various fields, e.g., reinforcement learning [58], domain generalization [24], multimodal large language models [4], Re-identification [27]. Recently, HQS [52] treats each label as task-specialized and employs the MoE model for learning in MLC tasks. However, unlike HQS, which assigns experts based on label-aware representation, our approach assigns experts based on group-aware representation, which can balance correlative (CO) and discriminative (DC) relationships to eliminate the risk of overfitting.
Figure 2: The overview of our framework. We apply clustering strategies to the co-occurrence graph to group labels into co-occurrence groups (CO) and discriminative groups (DC). Then, we introduce two types of prompt tokens into ViT, corresponding to CO and DC, respectively. ViT subsequently generates two types of group-aware representations. To adaptively map these representations to label-aware ones, we design Group-Aware MoEs. Finally, two distinct classifiers are employed to classify the label-aware representations.
3 Preliminaries
3.1 Notations
Let $\bm{x}\in\mathcal{X}$ denote the instance, and $\bm{y}\in\mathcal{Y}$ denote the corresponding label, where $\mathcal{X}=\mathbb{R}^{d}$ is the instance space, and $\mathcal{Y}=\{0,1\}^{K}$ is the label space containing the $K$ classes. For a given instance $\bm{x}$, $y^{k}=1$ denotes $k$-th label is a relevant label associated with the instance $\bm{x}$, and vice versa. In MLC tasks, given a labeled dataset with $N$ samples $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}$. Our goal is to train a model, denoted as $f(\bm{x};\theta):\mathcal{X}\rightarrow\mathcal{Y}$, using a training dataset $\mathcal{D}$. The aim is to accurately predict all potential labels $\bm{y}$ for unseen instances $\bm{x}$, with $\theta$ representing the learnable model parameters. Accordingly, the MLC expected risk function:
$\mathcal{R}\left(f\right)={}$ $\underset{\left(\bm{x}\!,\bm{y}\right)\sim p\left(\bm{x}\!,\bm{y}\right)}{\mathbb{E}}\left[\mathcal{L}\left(f\left(\bm{x}\right),\bm{y}\right)\right]~{}\text{,}$ (1)
where $\mathcal{L}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ indicates a multi-label loss function that turns MLC task into multiple more straightforward binary classification problems.
3.2 Visual Prompt Tuning Revisit
Visual Prompt Tuning is mainly implemented in ViTs [12], which usually consists of a patch embedding layer and $L$ stacked transformer blocks (Block). Given an image $\bm{x}$ that is divided into a set of non-overlapping patches, the patch embedding layer embedded each patch with the position information $\mathbf{E}_{\text{pos}}$ into a $D$-dimensional space, as follows:
$\mathbf{E}_{0}={\rm PatchEmbed}(\mathbf{x})+\mathbf{E}_{\text{pos}}~{}\text{.}$ (2)
Next, image patch embeddings $\mathbf{E}_{i}=\{\mathbf{e}_{i}^{j}\in\mathbb{R}^{D}|j\in\mathbb{N},1\leq j\leq N_{e}\}$ along with a class token $\mathrm{cls}_{i}$ is input into the ($i$+1)-th transformer block:
$[\mathrm{cls}_{i},\mathbf{E}_{i}]={\rm Block}_{i}([\mathrm{cls}_{i-1},\mathbf{E}_{i-1}])~{}\text{,}$ (3)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicates stacking and concatenation on the sequence length dimension.
For VPT, Jia et al. [21] advocate introducing a set of N learnable prompt tokens as input to $i$-th transformer block (${\rm Block}_{i}$), denoted as $\mathbf{P}_{i-1}=\{\mathbf{p}_{i-1,k}\in\mathbb{R}^{D}|k\in\mathbb{N},1\leq k\leq N_{p}\}$. The whole VPT is formulated as follows:
$[\mathrm{cls}_{i},\_,\mathbf{E}_{i}]\!=\!{\rm Block}_{i}([\mathrm{cls}_{i-1},\mathbf{P}_{i-1},\mathbf{E}_{i-1}])~{}\text{.}$ (4)
3.3 How to Group Labels?
To group the labels, we take an approach [50], which performs spectral clustering on the co-occurrence graph $\mathcal{G}^{+}$ and dis-occurrence graph $\mathcal{G}^{-}$. Specifically, we define a co-occurrence matrix $\mathbf{S}\in\mathbb{R}^{K\times K}$ and an affinity matrix,
[…]
where $\tau$ is a hyper-parameter. By applying spectral clustering to $\mathbf{M}$, we divide the classes $\mathcal{C}$ into several subsets within both CO and DC, designated as $\{\mathcal{C}^{+}_{t}\}^{N_{c}}_{t=1}$ and $\{\mathcal{C}^{-}_{t}\}^{N_{c}}_{t=1}$, respectively, where $N_{c}$ is the number of subsets in each set. Additionally, the number of classes $|\mathcal{C}^{+}_{t}|$/$|\mathcal{C}^{-}_{t}|$ within these subsets is expected to vary. More can be found in Appendix.
4 Method
In this section, we describe our novel framework, named ML-VPT, with group-aware visual prompt tuning (in §4.2) and group-aware MoE (in §4.3) for MLC, depicting the overall overview in Fig. 2. In the end, we describe in §4.4 how our CO and DC jointly optimize and infer.
4.1 Overview
To address the co-occurrence overfitting problem as well as high resource demands in training, we propose a novel multi-label learning method with VPT, and the key is to balance CO and DC, effectively reducing the overfitting risk resulting from overemphasized co-occurrence. To this end, first, we divide all classes into $N_{c}$ CO groups and $N_{c}$ DC groups as class subsets, according to the co-occurrence graph.
Secondly, to model CO and DC between labels, we add two sets of prompt tokens inside ViT, called CO and DC tokens $\mathbf{P}^{+}$/$\mathbf{P}^{-}$, by leveraging the capabilities of advanced VPT. In each subset, these tokens are evenly distributed across several groups, with each group associated with a class subset. Within the ViT, extended prompt tokens learn CO and DC relationships among labels, aiming to mitigate the adverse effects of CO relationships. Upon completing those, the ViT generates multiple group-aware representations $\mathbf{Z}^{+}$/$\mathbf{Z}^{-}$ for class subsets in CO and DC groups:
$\mathbf{Z}^{+}\cup\mathbf{Z}^{-}={\rm ViT}_{\Phi}([\mathbf{P}^{+},\mathbf{P}^{-}],\bm{x})~{}\text{.}$ (6)
These representations are shared among various classes within the same grouping strategy, which complicates the representation from group-aware to label-aware.
Subsequently, to enhance the learning of label-aware representations $\mathbf{C}^{+}$/$\mathbf{C}^{-}$ from group-aware representations for each class, we use MoEs to adaptively map these representations from group-aware to label-aware:
$\mathbf{C}^{+}\!\!=\!{\rm MoE}\biguplus{\rm Gate}(\mathbf{Z}^{+})\text{,}~{}~{}\mathbf{C}^{-}\!\!=\!{\rm MoE}\biguplus{\rm Gate}(\mathbf{Z}^{-})~{}\text{,}$ (7)
[…]
Finally, we employ two sets of classifiers with non-shared parameters, each making predictions in their respective, and then combine the results using a weighted sum.
4.2 Group-Aware Visual Prompt Tuning
Learning the relationship between labels has long been a key focus in MLC [7, 30]. However, overemphasizing these correlations often results in inferring objects primarily from co-occurring ones, which is unsuitable for all real-world scenarios [50]. On the other hand, learning discriminative representations for each object independently can lead to incorrect inferences when context or co-occurring objects are absent [46]. To address this issue, we decompose label relationships into two aspects: In the CO, learning shared information for each group is utilized to model the relationships between co-occurring labels. Conversely, in the DC, learning unique information for each group is employed to define the relationships between labels that do not co-occur.
Specifically, we define two types of prompt tokens, referred to as CO group prompts $\{\mathbf{p}^{+}_{t}\}^{N_{c}}_{t=1}$ and DC group prompts $\{\mathbf{p}^{-}_{t}\}^{N_{c}}_{t=1}$. On the one hand, from the perspective of grouping class, each prompt token $\mathbf{p}^{+}_{t}$/$\mathbf{p}^{-}_{t}$ uniquely maps to a specific class subset $\mathcal{C}^{+}_{t}$/$\mathcal{C}^{-}_{t}$. On the other hand, in terms of representation learning, the prompt token $\mathbf{p}^{+}_{t}$/$\mathbf{p}^{-}_{t}$ is associated with all classes, contained by $\mathcal{C}^{+}_{t}$/$\mathcal{C}^{-}_{t}$, thus forming a one-to-many relationship. For example, if $\mathcal{C}^{+}_{t}$ includes the classes {bicycle;car}, then $\mathbf{p}^{+}_{t}$ is exclusively associated with these classes, which can indicate vision representations of these classes. Similarly, $\mathbf{p}^{-}_{t}$ corresponds to $\mathcal{C}^{-}_{t}$. In the context of prompt tuning, we introduce a set of CO group prompts $\mathbf{P}_{i-1}^{+}=\{\mathbf{p}_{i-1,t}^{+}\in\mathbb{R}^{D}|t\in\mathbb{N},1\leq t\leq N_{c}\}$ and a set of DC group prompts $\mathbf{P}_{i-1}^{-}=\{\mathbf{p}_{i-1,t}^{-}\in\mathbb{R}^{D}|t\in\mathbb{N},1\leq t\leq N_{c}\}$ into each transformer block (${\rm Block}_{i}$), which is designed to facilitate the learning of visual representations pertinent to specific class subsets $\mathcal{C}^{+}_{t}$ and $\mathcal{C}^{-}_{t}$ [^1] The index $t$, in class subsets $\mathcal{C}_{t}^{+}$ and $\mathcal{C}_{t}^{-}$, may not be the same. However, for simplicity, we will uniformly denote it as $t$ in our analysis.. It is expressed in the following mathematical form:
$[\mathrm{cls}_{i},\_,\_,\mathbf{E}_{i}]={\rm Block}_{i}([\mathrm{cls}_{i-1},\mathbf{P}_{i-1}^{+},\mathbf{P}_{i-1}^{-},\mathbf{E}_{i-1}])~{}\text{.}$ (8)
Leveraging the powerful long-range modeling advantages of the transformer, each $\mathbf{p}^{+}_{i,t}$/$\mathbf{p}^{+}_{i,t}$ can aggregate semantic information about the class subset $\mathcal{C}_{t}$ from the image embedding $\mathbf{E}_{i}$. In CO, $\mathbf{p}_{i,t}^{+}$ focuses on learning shared information within the class subset $\mathcal{C}_{t}^{+}$. Conversely, in DC, due to the classes that do not co-occur in subsets $\mathcal{C}_{t}^{-}$, the DC group prompts $\mathbf{p}_{i,t}^{-}$ seek to capture the distinct information unique to each class within the subset. However, due to MSA, prompt tokens $\mathbf{p}_{i,t}$ may transfer information to each other. To mitigate this effect, we introduce a predefined mask to prevent information exchange between prompt tokens across groups, making them more discriminative.
4.3 Group-Aware Mixture-of-Experts
As in §4.2, each class subset was associated with one prompt token. However, this simplification does not accurately represent the real-world scenarios. For example, in the CO, the classes {traffic light;stop sign;car} exhibit a strong co-occurrence relationship and are classified into the same group. However, when only one prompt token is used, it becomes challenging to capture the co-occurrence relationships between labels within the same group. In contrast, for the DC, the classes {skateboard;skis;refrigerator}, which are divided into the same group due to their solid discrimination relationships, tend to show significant visual differences [54] in the images. When only one prompt token is assigned to each group, the model tends to share the group’s common semantic information, which in turn impedes its ability to capture the unique representations for each class.
To this end, we expand the number of prompt tokens for each group. Since each class possesses unique characteristics, it is essential to select a group-aware representation that is appropriate for each class. To achieve this, we frame the selection of group-aware representations as a multi-task learning problem, using the MoEs to play its advantages in multi-task learning [24]. And each subtask is associated with multiple group-aware representations, which are then mapped to label-aware representations via the MoE model.
More Prompt Tokens for Each Group. Specifically, we extend each prompt token $\mathbf{p}_{i-1,t}$ from 1 to $N_{m}$, defined as follows: $\mathbf{p}_{i-1,t}=\{\mathbf{p}_{i-1,t}^{e}\in\mathbb{R}^{D}|e\in\mathbb{N},1\leq e\leq N_{m}\}$. Moreover, $\mathbf{p}_{i-1,t}^{e,+}$ indicates the correlative group prompt tokens, while $\mathbf{p}_{i-1,t}^{e,-}$ is the opposite. To reduce additional parameters, we advocate introducing MoE after the last transformer block. Then, the final transformer block’s output can be expressed as: $[\mathrm{cls}_{o},\mathbf{Z}^{+},\mathbf{Z}^{-},\mathbf{E}_{o}]$, where $\mathbf{Z}^{+}$ corresponds to the output of the CO and $\mathbf{Z}^{-}$ is the opposite. More specifically, $\mathbf{z}_{t}^{e,+}$ is related to the $e$-th prompt token in the $t$-th discriminative group. And, we refer to $\mathbf{z}_{t}^{e}$ as the $e$-th representation corresponding to the class subset $\mathcal{C}_{t}$.
Mixture-of-Experts. Subsequently, we introduce correlative and discriminative group experts $\mathcal{E}$, and each expert $E_{t}^{e}(\cdot)$ is implemented as an FFN with a ReLU activation and residual connection:
+=$\mathcal{E}^{+}=\{{E}_{t}^{e,+}|e,t\in\mathbb{N},1\leq e\leq N_{m},1\leq t\leq N_{c}\}~{}\text{,}$ (9)
=$\mathcal{E}^{-}=\{{E}_{t}^{e,-}|e,t\in\mathbb{N},1\leq e\leq N_{m},1\leq t\leq N_{c}\}~{}\text{,}$ (10)
$E_{t}^{e}(\mathbf{z}_{t}^{e})=\mathbf{z}_{t}^{e}\!+\!\sigma(\mathbf{z}_{t}^{e}\mathbf{W}_{t,{\rm dn}}^{e}\!+\!\mathbf{b}_{t,{\rm dn}}^{e})\mathbf{W}_{t,{\rm up}}^{e}\!+\!\mathbf{b}_{t,{\rm up}}^{e}~{}\text{,}$ (11)
where $\mathbf{W}_{t}^{e}$ and $\mathbf{b}_{t}^{e}$ are the weights and biases of the experts, and $\sigma(\cdot)$ is a activation. To make lightweightness, the hidden dimension is kept small, e.g., 5. The two groups of experts don’t share parameters.
Gating Networks. We employ two distinct strategies for expert selection: Each subset within the grouping class utilizes a routing network to learn varying weights for multiple group-aware representations $\mathbf{z}_{t}^{e}$, followed by weighted merging by multiple experts. The second strategy considers each class independently; here, different classes within the corresponding class subset group employ different dynamic routing networks to determine diverse weights for multiple visual representations $\mathbf{z}_{t}^{e}$, subsequently merging these using multiple experts. We opt for the second strategy, as different classes necessitate distinct information, rendering identical shared visual representations insufficient.
For each class $k$, we define a gating network $g^{+}_{k}(\cdot)$/$g^{-}_{k}(\cdot)$ in CO and DC groups, respectively, implemented as a simple fully connected network:
$\bm{w}_{k}^{+}$ $={\rm softmax}(\mathbf{z}_{t}^{e,+}\mathbf{W}_{k}^{+}+\mathbf{b}_{k}^{+})\in\mathbb{R}^{N_{m}}\text{,}~{}k\in\mathcal{C}_{t}^{+}~{}\text{,}$ (12)
$\bm{w}_{k}^{-}$ $={\rm softmax}(\mathbf{z}_{t}^{e,-}\mathbf{W}_{k}^{-}+\mathbf{b}_{k}^{-})\in\mathbb{R}^{N_{m}}\text{,}~{}k\in\mathcal{C}_{t}^{-}~{}\text{,}$ (13)
where $\bm{w}_{k}^{+}\!=\!\{w_{k}^{e,+}\}^{N_{e}}_{e=1}$ and $\bm{w}_{k}^{-}\!=\!\{w_{k}^{e,-}\}^{N_{e}}_{e=1}$ denote the weights associated with class $k$ in class subset $\mathcal{C}^{+}_{t}$ and $\mathcal{C}^{-}_{t}$.
Adaptive Label-Aware Representation. Adaptively aggregate group-aware representations into label-aware representations. The group-aware MoE is formalized as follows:
$\mathbf{c}_{k}^{+}$ $=\sum\nolimits_{e=1}^{N_{e}}{w}_{k}^{e,+}E_{t}^{e,+}(\mathbf{z}_{t}^{e,+})~{}\text{,}$ (14)
$\mathbf{c}_{k}^{-}$ $=\sum\nolimits_{e=1}^{N_{e}}{w}_{k}^{e,-}E_{t}^{e,-}(\mathbf{z}_{t}^{e,-})~{}\text{.}$ (15)
where $\mathbf{c}_{k}^{+}$ and $\mathbf{c}_{k}^{-}$ indicate the label-aware representations for class $k$, obtained from CO and DC groups, respectively.
4.4 Optimization and Inference
In this work, we adapt dual classification heads to mitigate the negative impacts arising from the accumulated prediction errors from a single classification head [49]. The logits for both correlative and discriminative groups can be predicted using two linear classifiers with a sigmoid as follows:
$\hat{{y}}_{k}^{+}=\sigma(\mathbf{W}_{k}^{\top}\mathbf{c}_{k}^{+}+{b}_{k}^{+})~{},\quad\hat{{y}}_{k}^{-}=\sigma(\mathbf{W}_{k}^{\top}\mathbf{c}_{k}^{-}+{b}_{k}^{-})~{}.$ (16)
Here, $\hat{{y}}_{k}^{+}$ and $\hat{{y}}_{k}^{-}$ represent the predictions for class $k$ in the two types of , while $f^{+}(\bm{x})=\hat{\bm{y}}^{+}$ and $f^{-}(\bm{x})=\hat{\bm{y}}^{-}$ denote the predictions for all classes about instance $\bm{x}$, respectively.
Referring to Eq. (1), the overall expected risk associated with our proposed method can be reformulated as follows:
$\mathcal{R}\!\left(f\right)\!=\!\!\underset{\left(\bm{x}\!,\bm{y}\right)\sim p\left(\bm{x}\!,\bm{y}\right)}{\mathbb{E}}\!\!\left[\mathcal{L}\left(f^{+}\left(\bm{x}\right),\bm{y}\right)\!\!+\!\mathcal{L}\left(f^{-}\left(\bm{x}\right),\bm{y}\right)\right]\text{,}$ (17)
where $\mathcal{L}(\cdot,\cdot)$ is ASL loss function [1]. Given an instance $\bm{x}$, the model’s final prediction is $f(\bm{x})=\hat{\bm{y}}=0.5\cdot(\hat{\bm{y}}^{+}+\hat{\bm{y}}^{-})$.
5 Experiments
Table 1: Comparison of our method with SOTA models on COCO at 224 $\times$ 224 and 448 $\times$ 448 resolution. All metrics are in %. Since mAP, CF1, and OF1 are among the most important evaluation metrics, they are highlighted in dark gray in our method.
                    Resolution: 224 $\times$ 224                 Resolution: 448 $\times$ 448
Method Backbone Pre-trained Data mAP CP CR CF1 OP OR OF1 mAP CP CR CF1 OP OR OF1
VPT 78.0 71.7 72.3 72.0 74.6 75.4 75.0 82.6 76.1 76.7 76.4 78.2 79.1 78.7
GateVPT 75.6 69.7 70.3 70.0 72.9 73.7 73.3 80.4 74.2 74.7 74.4 76.6 77.5 77.0
E2VPT 77.3 71.6 72.2 71.9 74.5 75.4 75.0 81.7 75.3 75.9 75.6 77.7 78.5 78.1
Ours ViT-B ImageNet 1K 79.6 73.1 73.6 73.3 75.7 76.5 76.1 83.6 76.9 77.5 77.2 79.1 79.9 79.5
VPT 71.0 66.0 66.6 66.3 71.1 72.0 71.6 72.2 66.9 67.6 67.3 72.3 73.1 72.7
GateVPT 66.5 62.4 62.9 62.6 68.2 69.0 68.6 69.1 64.5 65.1 64.8 70.1 70.9 70.5
E2VPT 69.8 65.0 65.6 65.3 70.4 71.2 70.8 73.0 67.8 68.4 68.1 72.8 73.6 73.2
Ours MAE ImageNet 1K 75.2 69.3 69.9 69.6 74.0 74.9 74.4 78.8 72.8 73.4 73.1 76.8 77.6 77.2
VPT 73.0 67.6 68.1 67.9 71.8 72.6 72.2 75.9 70.2 70.8 70.5 74.1 74.9 74.5
GateVPT 70.9 65.7 66.3 66.0 70.3 71.2 70.8 74.9 69.4 70.0 69.7 73.4 74.2 73.8
E2VPT 72.7 67.3 67.9 67.6 71.7 72.5 72.1 76.0 70.3 70.9 70.6 74.2 75.0 74.6
Ours MoCo v3 ImageNet 1K 75.1 69.1 69.7 69.4 73.2 74.0 73.6 77.3 71.4 72.0 71.7 75.2 76.0 75.6
VPT 81.0 74.5 75.1 74.8 77.3 78.1 77.7 85.0 78.4 79.0 78.7 80.5 81.4 81.0
GateVPT 80.8 74.2 74.8 74.5 76.9 77.8 77.3 84.5 77.8 78.4 78.1 80.1 80.9 80.5
E2VPT 81.9 75.4 76.0 75.7 77.9 78.8 78.3 85.2 78.6 79.2 78.9 80.7 81.6 81.1
Ours ViT-B-21k ImageNet 21K 83.0 76.2 76.7 76.4 78.6 79.5 79.0 86.4 79.7 80.3 80.0 81.6 82.5 82.0
VPT 86.1 79.5 80.1 79.8 81.9 82.8 82.4 89.7 83.2 83.8 83.5 85.2 86.1 85.7
GateVPT 85.6 79.0 79.6 79.3 81.6 82.5 82.0 89.1 82.5 83.1 82.8 84.5 85.5 85.0
E2VPT 86.3 79.8 80.4 80.1 82.2 83.1 82.6 89.6 83.1 83.7 83.4 85.2 86.1 85.6
Ours DINOv2/B LVD-142M [35] 87.5 80.8 81.4 81.1 83.0 83.9 83.4 90.6 84.2 84.8 84.5 86.0 86.9 86.4
VPT 80.1 73.8 74.4 74.1 77.1 77.9 77.5 84.5 78.2 78.8 78.5 81.1 82.0 81.5
GateVPT 79.4 73.2 73.8 73.5 76.5 77.3 76.9 83.3 77.0 77.7 77.3 80.0 80.9 80.4
E2VPT 80.6 74.2 74.8 74.5 77.5 78.4 77.9 84.3 78.1 78.8 78.5 81.0 81.9 81.4
Ours DINOv2/S LVD-142M [35] 83.4 76.6 77.1 76.8 79.6 80.5 80.0 87.4 80.8 81.4 81.1 83.2 84.1 83.7
5.1 Experimental Settings
Dataset and Evaluation Metric. In this work, we evaluate the effectiveness of the proposed method on four benchmark datasets, including Pascal VOC 2007 (VOC07) [13], MS-COCO 2014 (COCO) [26], NUS-WIDE (NUS) [8], and Visual Genome (VG256) [23]. For a fair comparison, we utilize widely adopted metrics: Mean Average Precision (mAP) across all classes. Additionally, consistent with previous works [7, 33], we also showcase overall precision (OP), recall (OR), F1-measure (OF1), as well as per-category precision (CP), recall (CR), and F1-measure (CF1) for detailed comparisons. It is important to emphasize that mAP, OF1, and CF1 are the most important metrics among these.
Implementation Details. Following previous MLC methods [37, 30, 33], we adopt a similar experimental setup. The AdamW [32] optimizer with the one-cycle policy lr_schedule [40] is applied to train the model with maximal learning rate of 0.0005. All models are trained for 40 epochs with the early stopping. The batch size is set to 64. For data augmentation, we apply RandAugment [9] and Cutout [10]. To make the model more robust, we apply an exponential moving average to the model parameters $\theta$, using a decay rate of 0.9997.
Pre-trained Models. To verify the robustness of our proposed method, we utilized a series of pre-trained models as backbone networks, including ViT [12], MAE [18], MoCo v3 [6], and DINOv2 [35]. For ViT, we employ ViT-B and ViT-B-21k, supervised training on ImageNet 1k [39], and ImageNet 21k [39] respectively. MAE and MoCo v3 are trained on ImageNet 1k for autoregressive and contrastive self-supervised training, respectively. For DINO v2, we utilize DINOv2/B and DINOv2/S, the former is the base model and the latter is the small variant, trained on a larger dataset for self-supervised learning.
Comparing Methods. To evaluate the performance of our model, we carried out a comprehensive set of experiments involving the following methods: VPT [21], GateVPT [53], and E2VPT [17]. These methods are primarily designed for single-label tasks, and the $\mathrm{cls}$ token output of the model is input into a linear classifier and optimized using the ASL [1] to adapt it to multi-label tasks. In our proposed methods, by default, the number of groups in the CO and DC groups is 5, and the number of experts in each group is 3. In the work, all comparing methods are setup and compared fairly.
5.2 Compared to State-of-the-Art (SOTA) Results
Performance on MS-COCO 2014. In Tab. 1, we report the comparison results between our method and the state-of-the-art method on the COCO. This table demonstrates that, under the same setting (i.e., the same pre-trained backbone and resolution), our method generally surpasses SOTA methods in terms of mean mAP, CF1, and OF1 at both 224$\times$224 and 448$\times$448 resolutions. Specifically, at the 224$\times$224 resolution, our method achieves improvements over the SOTA methods as follows: the mAP increases by 1.2%-8.7%, the CF1 by 0.7%-7.0%, and the OF1 by 0.7%-5.8%. Similarly, at the 448$\times$448 resolution, improvements are noted in the mAP from 0.4% to 6.6%, the CF1 from 0.8% to 5.8%, and the OF1 score from 0.7% to 4.5%, also compared with the SOTA method. These results demonstrate the effectiveness of our method. Moreover, it should be noted that GateVPT and E2VPT, which are modified from VPT, do not consistently outperform VPT. For instance, at a 224$\times$224 resolution using ViT-B, MAE, and MoCo v3, VPT demonstrates superior performance compared to these two modifications. A potential explanation could be that GateVPT and E2VPT are tailored for single-label tasks and do not account for label correlations during the feature extraction phase. The best mAP is 90.6% achieved by our method, using DINOv2/B at a resolution of 448$\times$448, which can be attributed to the superior capabilities of DINOv2/B and higher resolution.
Table 2: Comparison of our method with SOTA models on VOC07 at 224 $\times$ 224 and 448 $\times$ 448 resolution. All metrics are in %.
            Resolution: 224         Resolution: 448
Method Backbone mAP CF1 OF1 mAP CF1 OF1
VPT 90.4 83.4 85.7 93.4 87.4 89.1
GateVPT 87.4 80.9 82.9 89.2 83.1 85.2
E2VPT 85.8 79.4 83.7 92.4 86.5 88.3
Ours ViT-B 92.9 86.4 88.1 94.3 88.6 90.1
VPT 80.3 74.1 79.2 82.8 75.9 81.6
GateVPT 85.1 75.4 80.5 86.1 79.2 84.2
E2VPT 83.3 76.7 81.5 84.6 78.2 83.2
Ours MAE 89.2 82.7 86.4 90.4 83.9 87.4
VPT 86.2 79.7 83.6 88.4 82.3 85.7
GateVPT 83.5 77.1 81.2 87.7 81.1 84.7
E2VPT 85.9 79.2 83.1 88.8 82.4 85.6
Ours MoCo v3 91.1 84.2 86.8 92.7 86.3 88.6
VPT 94.2 88.0 89.6 94.4 88.7 90.4
GateVPT 94.3 88.5 89.7 95.2 89.6 91.1
E2VPT 94.1 87.9 89.4 95.1 89.3 90.7
Ours ViT-B-21k 95.0 89.2 90.3 95.6 90.0 91.3
VPT 95.6 90.2 91.6 96.1 91.0 92.3
GateVPT 95.5 89.7 91.2 96.2 91.1 92.5
E2VPT 95.4 89.9 91.4 96.1 91.1 92.4
Ours DINOv2/B 96.4 91.0 92.3 97.0 92.1 93.1
VPT 92.8 86.4 88.5 93.8 88.1 90.1
GateVPT 92.1 85.5 87.7 93.3 87.3 89.3
E2VPT 89.6 83.1 86.3 91.2 84.8 88.0
Ours DINOv2/S 94.3 88.3 90.3 95.7 90.2 92.0
Performance on Pascal VOC 2007. For VOC07, we present the results of both our method and the comparison method in Tab. 2. For clarity and brevity, unlike previous methods [7, 30, 42], we report the same evaluation metric as those used in the COCO and do not provide the Average Precision (AP) for each class. Table 2 illustrates that in the same settings, our method generally outperforms other approaches in mAP, OF1, and CF1 at both 224$\times$224 and 448$\times$448 resolutions. Specifically, at the 224$\times$224 resolution, our method shows improvements of 0.7%-8.9% in mAP, 0.7%-8.6% in CF1, and 0.6%-7.2% in OF1 compared to SOTA methods. Similarly, at the 448$\times$448 resolution, the enhancements are 0.4%-7.6% in mAP, 0.4%-8.0% in CF1, and 0.2%-5.8% in OF1 compared to SOTA methods. These results robustly confirm the effectiveness of the proposed method.
Table 3: Comparison of our method with SOTA models on NUS and VG256 at 224 $\times$ 224 resolution. All metrics are in %.
             NUS          VG256
Method Backbone mAP CF1 OF1 mAP CF1 OF1
VPT 65.2 62.8 73.8 42.2 44.0 57.5
GateVPT 62.8 61.5 72.9 41.1 43.0 56.5
E2VPT 63.9 62.2 73.7 42.2 44.2 57.6
Ours ViT-B 65.7 63.3 74.1 44.2 45.8 58.8
VPT 60.3 59.3 73.0 37.9 40.3 55.2
GateVPT 58.2 57.6 71.9 35.3 38.2 53.2
E2VPT 60.6 59.4 73.0 37.6 40.2 55.0
Ours MAE 61.9 60.7 73.5 40.5 42.7 57.3
VPT 62.7 61.3 73.8 40.4 42.6 56.5
GateVPT 61.8 60.5 72.9 39.6 41.8 55.8
E2VPT 62.8 61.2 73.5 40.4 42.5 56.5
Ours MoCo v3 63.0 61.1 73.6 41.0 43.1 57.1
VPT 67.5 64.8 74.7 46.0 47.4 60.2
GateVPT 66.9 64.4 74.3 45.4 46.8 59.5
E2VPT 67.7 64.9 74.7 46.0 47.4 60.1
Ours ViT-B-21k 68.2 65.0 75.0 47.6 48.6 61.2
VPT 68.1 65.1 75.0 49.3 50.3 62.9
GateVPT 67.2 64.6 74.5 48.6 49.7 62.4
E2VPT 67.9 65.0 75.0 49.2 50.3 62.9
Ours DINOv2/B 68.7 65.5 75.2 50.9 51.7 64.1
VPT 65.3 63.2 74.3 45.2 46.7 60.2
GateVPT 64.5 62.6 73.7 44.3 45.9 59.3
E2VPT 65.4 63.1 74.3 45.1 46.7 60.1
Ours DINOv2/S 66.9 64.3 74.8 47.7 48.7 61.9
Performance on NUS-WIDE and Visual Genome. In Tab. 3, we show the experimental results for NUS and VG256. For the VG256 setup, we follow previous work [49]. Our method achieves the best performance and outperforms SOTA methods in mAP, CF1, and OF1, under the same pre-training model and resolution setting.
5.3 Diagnostic Experiments
Impact of Different Components. To investigate the impact of different components of our method, including Group-Level Visual Prompt Tuning (GVPT) and Group-Level Mixture-of-Experts (GMoE), we conducted experiments on two benchmark datasets: COCO and VOC07. The Fig. 3 illustrates the performance enhancements achieved by GVPT on the COCO, with improvements of 1.41% in mAP, 1.22% in CF1, and 1.1% in OF1, respectively. Moreover, incorporating GMoE into GVPT further improves performance by 0.99%, 0.72%, and 0.49%, respectively, compared to GVPT alone. On the VOC07, GVPT led to improvements of 1.77%, 1.97%, and 1.55% in mAP, CF1, and OF1, respectively. However, GMoE has limited improvement in model performance, e.g., 0.48% in mAP. These results reveal that the grouping classes in VOC07 have a relatively weak ability to learn label correlations within each group, as each group contains relatively few classes, and the distinctions between multiple tokens within each group may be minimal.
(a) evaluation on COCO
(b) evaluation on VOC07
Figure 3: The analysis of different components. VPT stands for the vanilla model, GVPT&GMoE combines both GVPT and GMoE. Note that each model uses 30 additional prompt tokens.
Grouping Strategies. To study the impact of the grouping strategy, we use four settings: vanilla VPT, using only the CO group, using only the DC group, and using both. As can be seen from Fig. 4, on COCO and VOC07, using the CO group and the DC group, respectively, has a certain improvement. When using both modes, the model is improved even more. We analyze that emphasizing correlative and discriminative relationships among labels has a certain gain on the MLC. While emphasizing the co-occurrence relationship, cleverly using the discriminative relationship will further compete for the model’s performance.
(a) evaluation on COCO
(b) evaluation on VOC07
Figure 4: The analysis of different grouping strategies. VPT is vanilla mode; VPT-CO indicates the VPT uses the CO groups. VPT-DC applies DC groups in VPT. GVPT incorporates both. Note that each model uses 10 additional prompt tokens and excludes MoE.
Number of Groups. In this study, our main idea is to group classes within the co-occurrence graph. We then aim to establish both a CO and DC relationship. Hence, additional research is required to explore the impact of group quantity on model performance. As illustrated in Fig. 5, we present comprehensive ablation studies on the COCO and VOC07. For the COCO, model performance improves with the increased number of groups when the same pre-trained model is used. Notably, when pre-training with MoCo v3, optimal performance is achieved with 5 groups. For VOC07, optimal results are achieved when the number of groups is set to either 5 or 8. The potential reason is the dataset’s relatively small number of 20 classes, which likely impedes the effective learning of label correlations in the CO group. Additionally, the average number of labels per image in VOC07 is only 1.5, leading to weaker correlations between labels in many images compared to those in COCO.
(a) mAP on COCO
(b) mAP on VOC07
Figure 5: The performance curve varies with the increase in the number of groups. More can be found in Appendix.
Number of Experts. The number of experts corresponds to the number of group-level representatives within each class group, which reflects the capacity to model label relationships within each group. Accordingly, we study the influence of the number of experts on model performance. As shown in Fig. 6, on the COCO, model performance marginally improves as the number of experts increases. However, the performance on VOC07 shows slight fluctuations. This phenomenon may be attributed to the small number of classes in each group, which does not require more prompt tokens to learn the relationship between labels. Therefore, adding more experts yields limited benefits and could potentially detract from model performance.
(a) mAP on COCO
(b) mAP on VOC07
Figure 6: The performance curve varies with the increase in the number of experts, in the mAP, More can be found in Appendix.
Table 4: Computation cost and parameters number on COCO.
Method # Total param. # Learnable param. FLOPs mAP
Linear Probing 86.64 M 0.06 M 21.96 G 77.8%
Fine Tuning 86.64 M 86.64 M 21.96 G 84.2%
VPT 86.91 M 0.33 M 24.51 G 86.1%
E2VPT 87.01 M 0.43 M 24.55 G 86.3%
GateVPT 86.67 M 0.08 M 24.51 G 85.6%
GVPT (Ours) 86.98 M 0.39 M 24.51 G 87.2%
ML-VPT (Ours) 87.60 M 1.02 M 24.51 G 87.5%
Figure 7: Visualization of model predictions on COCO. Each example includes an input image, the co-occurrence probability, and the predictions of the two models.
Computation Cost and Parameters. To verify the first issue mentioned in §1, we use DINOv2/B as a pre-trained model in several methods to compare the computational overhead and parameters. As shown in Tab. 4, FineTuning not only requires more learnable parameters but also reduces model performance. Our method achieves the highest result with limited computational resources and parameters.
5.4 Case Study
To verify that our method can alleviate the co-occurrence overfitting problem, we select label pairs with high co-occurrence probabilities to study. As shown in Fig. 7, ML-VPT is compared with VPT on COCO. Under the influence of high co-occurrence probabilities, VPT is prone to over-inference, resulting in model misprediction. For example, in the second example, when toilet appears, VPT infers that there is a sink through co-occurrence classes and other contextual information (the sink has been removed). However, our method can correctly predict all labels. Similarly, in outdoor scenes, traffic light and stop sign are both misclassified as car, while in animal-related classes, the presence of a cow leads to the false inference of a person. These examples empirically demonstrate that ML-VPT can mitigate co-occurrence overfitting.
6 Conclusion
In MLC, mainstream methods tend to learn co-occurrence relationships between labels. However, an overemphasis on these relationships can lead to model overfitting. To address this issue, we introduce a novel Multi-Label Visual Prompt Tuning method designed to balance both correlative and discriminative relationships among labels. Our method offers two advantages: i) We employ the concept of grouping classes to independently model the correlative and discriminative relationships between labels, thus reducing the risk of overfitting. ii) While the grouping idea allows multiple classes to share the same group information, we further enhance our approach by integrating a group-aware MoE model that dynamically maps optimal multiple group-aware representations to a label-aware representation for each class.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (No.62076005, No.62472004, No.61876002), the Provincial Quality Project of Education in the New Era in 2023 (Postgraduate Education No.2023lhpysfjd009) and the University Synergy Innovation Program of Anhui Province, China (GXXT-2021-002, GXXT-2022-029, GXXT-2022-031). We thank the reviewers for their comments and suggestions. We also thank the High-performance Computing Platform of Anhui University for providing computational resources for this project.